Add one set of parentheses to the expression:
7*X+3+8*X+3*X+8-9 so that it is equivalent to 2(9X+10).

Answer by Edwin McCravy(20081)

7*(X+3)+8*X+3*X+8-9 =

7*(X+3)+11*X-1 =

7*X+21+11*X-1 =

18*X+20 =

2(9X+10)

Edwin